# Language Integration in Fine-Tuning Multimodal Large Language Models for Image-Based Regression

### v1

@@ 4.5 Generalization to AI-generated images
Table 6: Prompt-Gated Regression on AGIQA-3k. Models fine-tuned on a single task (rows 1 and 2), compared to a model fine-tuned on both tasks (row 3). All models trained with textual prompts. Parenthesis shows difference from baseline.
@@
Alignment 0.810 0.889 0.709 (-0.163) 0.793 (-0.123)
Perception 0.676 (-0.134) 0.781 (-0.108) 0.872 0.916
Alignment + Perception 0.804(-0.06) 0.885 (-0.04) 0.875 (+0.03) 0.913 (-0.03)
@@
Task-specific prompt sensitivity. In contrast, the perceptual quality task shows minimal sensitivity to prompt variations. Performance remains largely stable across all prompt conditions (SRCC: 0.872$\pm$0.006), indicating that visual quality assessment relies primarily on image features rather than semantic context. This differential sensitivity validates that prompt utility depends on task semantics: prompts enhance performance when semantically relevant but provide little benefit for purely visual assessment tasks.
@@ 5 Conclusions
This work challenges how multimodal large language models should be applied to image-based regression tasks. Our analysis reveals that effective multimodal regression requires moving beyond human-mimicking approaches. Three key insights emerge from our study: First, vocabulary constraints hinder rather than help regression performance; simple bin-based classification outperforms complex vocabulary-dependent methods. Second, semantic relevance in prompts is crucial for unlocking cross-modal capabilities, while generic task descriptions provide no meaningful benefit over image-only training. Third, architectural simplicity often trumps complexity; i.e. increasing classification bins vs. distributional modeling approaches. While our experiments focus on image assessment domains, the generalizability of these principles to other multimodal regression tasks remains an important area for future investigation. This work provides a foundation for developing more effective multimodal regression systems and underscores the critical role of semantic coherence in cross-modal understanding.
This work opens several research directions. First, investigating whether image-only fine-tuning protocols can preserve multimodal prompt-leveraging capabilities would enable more flexible and resource-efficient model development. Second, exploring multitask setups where regression and text generation tasks complement each other could lead to more capable unified models.

### v3

@@ 4.5 Generalization to AI-generated images
Table 8: Prompt-Gated Regression on AGIQA-3k. Models fine-tuned on a single task (rows 1 and 2), compared to a model fine-tuned on both tasks (row 3). All models trained with textual prompts.
@@
Alignment 0.810 0.889 0.709 0.793
Perception 0.676 0.781 0.872 0.916
Align. + Percep. 0.804 0.885 0.875 0.913
@@
[…]
@@ 5 Conclusions
This work challenges current multimodal large language model approaches to image-based regression tasks. Three key insights emerge: First, vocabulary constraints hinder performance; simple bin-based classification outperforms complex vocabulary-dependent methods. Second, semantic relevance in prompts is crucial for unlocking cross-modal capabilities, while generic task descriptions provide no benefit over image-only training. Third, simple bin-based classification often outperforms complex modeling approaches. Results across different backbones demonstrate that these principles generalize across various MLLM architectures. This work provides a foundation for more effective multimodal regression systems and underscores the critical role of semantic coherence in cross-modal understanding.
Future work. Future directions include evaluating RvTC beyond image assessment domains, exploring prompt generation for datasets without textual annotations, investigating whether image-only fine-tuning can preserve multimodal capabilities, and exploring multitask setups combining regression and text generation tasks.